# revtex4-2Repair the float

### v3

@@ 2 Landau-Ginzburg Theory
In this section we review the Landau-Ginzburg (LG) theory for a unidirectional PDW state [25], which will be used to describe the profiles of our order parameters in the presence of static topological defects. In a a later section we will calculate the Local-Density-of-States (LDOS) in the presence of topological defects that we will describe in this section. Because our defects are static, we will be neglecting fluctuations, so our analysis here will be done at the level of mean field theory. Furthermore, we also work deep in the PDW phase where the order parameters have well defined local amplitudes, which is justified at low temperatures.
@@ 2.1 Free energy of the PDW state
The free energy of the LG theory is chooses to have the form:
$\mathcal{F}[\Delta_{0},\Delta_{\mathbf{Q}},\Delta_{-\mathbf{Q}}]=\mathcal{F}_{\text{sc}}[\Delta_{0}]$ $+\mathcal{F}_{\text{pdw}}[\Delta_{\mathbf{Q}},\Delta_{-\mathbf{Q}}\big{]}$ (2.1)
The first term in equation (2.1) describes uniform superconductivity, which takes the familiar form
$\mathcal{F}_{\text{sc}}=\frac{1}{2m^{*}}\Big{|}\big{(}-i\nabla+2e\mathbf{A}\big{)}\Delta_{0}\Big{|}^{2}+\frac{a}{2}|\Delta_{0}|^{2}+\frac{b}{2}|\Delta_{0}|^{4}$ (2.2)
@@
The second term in equation (2.1) describes the free energy of the unidirectional PDW. In Eq.(1.1) we presented the expansion of the local pairing amplitude $\Delta(\bm{r})$ in its Fourier components $\Delta_{\pm\bm{Q}}(\bm{r})$, where $\Delta_{\pm\bm{Q}}(\bm{r})$ are the order parameters of the unidirectional PDW state with ordering wave vector $\bm{Q}$. Since the local pairing amplitude $\Delta(\bm{r})$ is a complex field, the two PDW components $\Delta_{\pm\bm{Q}}$ are not the complex conjugate of each other (as they would have been for a CDW) but are two independent complex fields. The PDW free energy has the form [25, 26, 45, 1]:
$\mathcal{F}_{\text{pdw}}[\Delta_{\mathbf{Q}},\Delta_{-\mathbf{Q}}\big{]}=$ $\kappa\Big{(}|\nabla\Delta_{\mathbf{Q}}|^{2}+|\nabla\Delta_{-\mathbf{Q}}|^{2}\Big{)}+r\Big{(}|\Delta_{\mathbf{Q}}|^{2}+|\Delta_{-\mathbf{Q}}|^{2}\Big{)}+u\Big{(}|\Delta_{\mathbf{Q}}|^{2}+|\Delta_{-\mathbf{Q}}|^{2}\Big{)}^{2}+\gamma|\Delta_{\bm{Q}}|^{2}|\Delta_{-\bm{Q}}|^{2}$ (2.3)
where $r=T-T_{c}^{\rm pdw}$, and $T_{c}^{\rm pdw}$ is the (mean-field) critical temperature for the PDW superconductor. We will assume that the coupling constant $\gamma<0$ as needed for an attractive interaction needed for a PDW, an LO-type state. Instead, a repulsive value of the coupling constant, $\gamma>0$, favors an FF-type state.
@@
In many superconductors of interest, such as cuprates Bi2Sr2CaCu2O8+δ and La2-xBaxCuO4 in the Meissner state, a uniform superconducting order parameter $\Delta_{0}$ is present. This results in the following additional “lock-in” terms that couple the three superconducting order parameters and it must be added to the free energy of Eq.(2.1)
@@ 2.2 Order parameters of the PDW state
$\rho_{\bm{Q}}(\bm{r})$ $=\Delta_{0}(\bm{r})\Delta_{\bm{Q}}^{*}(\bm{r})$ (2.5)
$\rho_{2\bm{Q}}(\bm{r})$ $=\Delta_{\bm{Q}}(\bm{r})\Delta_{-\bm{Q}}^{*}(\bm{r})$ (2.6)
$\Delta_{\text{4e}}({\bm{r}})$ $=\Delta_{\mathbf{Q}}({\bm{r}})\Delta_{-\mathbf{Q}}({\bm{r}})$ (2.7)
@@
Following the analysis of Berg et al [25] we decompose the phase fields of the PDW order parameters as
$\theta_{\pm\bm{Q}}(\bm{r})=\theta_{+}(\bm{r})\pm\theta_{-}(\bm{r})$ (2.8)
@@
$\Delta_{\pm\bm{Q}}(\bm{r})$ $\to\exp(i(\vartheta_{+}\pm\vartheta_{-}))\Delta_{\pm\bm{Q}}(\bm{r})$ (2.9)
$\Delta_{0}(\bm{r})$ $\to\exp(i\vartheta_{0})\Delta_{0}(\bm{r})$ (2.10)
$\rho_{\bm{Q}}(\bm{r})$ $\to\exp(i(\vartheta_{0}-\vartheta_{+}))\exp(i\vartheta_{-})\rho_{\bm{Q}}(\bm{r})$ (2.11)
$\rho_{2\bm{Q}}(\bm{r})$ $\to\exp(i2\vartheta_{-})\rho_{2\bm{Q}}(\bm{r})$ (2.12)
$\Delta_{4e}(\bm{r})$ $\to\exp(i2\vartheta_{+})\Delta_{4e}(\bm{r})$ (2.13)
@@
$\rho(\bm{r})=\bar{\rho}+\rho_{\bm{K}}(\bm{r})\exp(i\bm{K}\cdot\bm{r})+\rho_{-\bm{K}}(\bm{r})\exp(-i\bm{K}\cdot\bm{r})+\ldots$ (2.14)
@@
$\mathcal{F}_{PDW-CDW}=g\,\rho_{\bm{K}}\Delta_{\bm{Q}}\Delta^{*}_{-\bm{Q}}+{\rm c.c.}$ (2.15)
where $g$ is a coupling constant [25]. Translation invariance then requires that the mutual commensurability condition is satisfied. The same requirement exists for a coupling between a spin density wave $SDW$. This effect is seen in La2-xBaxCuO4 at the charge-ordering transition [50] and in La1.8-zEu0.2SrxCuO4 [51] at low temperatures. The existence of this interaction yields some interesting physics not allowed for an uniform SC state: the PDW is sensitive to charged impurities due to their coupling to the CDW. Thus, unlike uniform superconducting order, which is only weakened by disorder, even small amounts of charge disorder destroys true long-range incommensurate PDW order.
@@ 2.3 Topological defects of the PDW state
$q_{v}=\frac{1}{2}(m_{\bm{Q}}+m_{-\bm{Q}}),\qquad q_{d}=\frac{1}{2}(m_{\bm{Q}}-m_{-\bm{Q}})$ (2.16)
We will label the topological defects by the combinations $(q_{v},q_{d})$, the vorticity and dislocation charges, or equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})$. In the simplest cases they are a) the superconducting (Abrikosov) vortex with topological charges $(q_{v},q_{d})=(\pm 1,0)$ (or, equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})=(\pm 1,\pm 1)$), b) the half-vortex (bound to a single CDW dislocation) with topological charges $(q_{v},q_{d})=(\pm 1/2,\pm 1/2)$ (or, equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})=(\pm 1,0)$), and c) the CDW double dislocation with topological charges $(q_{v},q_{d})=(0,\mp 1)$ (or, equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})=(\pm 1,\mp 1)$). These identifications imply that a conventional superconducting vortex is equivalent to both PDW order parameters $\Delta_{\pm\bm{Q}}$ having a vortex. The half-vortex is equivalent to a vortex in $\Delta_{\bm{Q}}$ but not in $\Delta_{-\bm{Q}}$ (and vice-versa) and it has a single-dislocation, as required by Eq.(2.6). Finally, a double dislocation is a vortex in $\Delta_{\bm{Q}}$ and an anti-vortex in $\Delta_{-\bm{Q}}$ (and vice-versa).
@@
In this paper will focus primarily on the properties of the half-vortex which we will regard as a static topological defect of an ordered PDW state which we will take to be of the Larkin-Ovchinnikov (LO) type [52]. Hence, we will assume that at long distances the PDW amplitudes are equal and constant in space, $|\Delta_{\bm{Q}}|=|\Delta_{-\bm{Q}}|$. However, at the core of the half-vortex one of these two amplitudes, say $\Delta_{\bm{Q}}$, must vanish while the other amplitude does not. As a result, the core of the half-vortex is in a Fulde-Ferrell (FF) state [24], and inversion symmetry is broken at the core of the half-vortex since $|\Delta_{\bm{Q}}|\neq|\Delta_{-\bm{Q}}|$. On the other hand, none of these considerations apply to either the Abrikosov vortex or to the double dislocation.
As in all superconductors, the Abrikosov vortex of the PDW arises in the presence of a magnetic field. On the other hand, the half-vortex and the double dislocation can appear due to the interaction of the PDW state with sufficiently strong charged impurities. This is possible since the CDW order parameter $\rho_{2\bm{Q}}(\bm{r})$ of the PDW has a linear coupling to charged impurities potentials whereas gauge invariance requires that the superconducting order parameters only interact quadratically through $|\Delta_{\pm\bm{Q}}(\bm{r})|^{2}$.
The properties of all three topological defects will be discussed in section 4, where we specify the form of the order parameters used in the numerics. The associated charge density wave patterns of each defect will be compared in section 4.2 as well as their spectral functions in 4.5.
An extremely useful experimental technique for detecting and visualizing CDW patterns is Scanning-Tunneling-Microscopy (STM) [53, 54, 55, 56]. STM has been used to investigate in great detail the charge order present in the superconducting phase of Bi2Sr2CaCu2O8+δ [57, 58, 59] and in the vortex halo [15, 16]. Relevant to the existence of PDW order is the experimental evidence for static (pinned) half-vortices in the superconducting phase of Bi2Sr2CaCu2O8+δ found in STM experiments by Du et al[14]. These authors argued that the jumps in the PDW SC winding-number are located around the charge dislocations. In this regime the term proportional to $\beta_{2}$ in Eq.(2.4) locks the phase of the PDW order parameter to the phase of the uniform superconducting state. The result is a confinement of the half-vortices into (half) vortex-anti-vortex pairs. Hence, in the phase in which the uniform order parameter is present, $\Delta_{0}\neq 0$, the energy of the half-vortex is linearly divergent and half-vortices cannot exist in isolation. In this regime charged impurity potentials can separate the half-vortices and anti-half-vortices as static defects.
@@ 3 The Half-Vortex
In this section we will model the profile of our half-vortex of an unidirectional PDW and we will assume that there is no uniform SC order. In section 4 we will embed the configuration of the half-vortex in a Bogoliubov-de Gennes Hamiltonian for the d-wave superconducting state of a CuO2 plane and investigate its effects on the electronic structure. The Abrikosov vortex and the double dislocation will also be considered in later sections, so we mention their solutions here as well. We should note that the associated CDW dislocation was given in Ref. [45].
We will seek an extremal solution of the Landau-Ginzburg free energy density $\mathcal{F}_{pdw}(\Delta_{\bm{Q}}(\bm{r}),\Delta_{-\bm{Q}}(\bm{r}))$ for the PDW order given in Eq.(2.3). In section 2 we showed that a vortex with half magnetic flux quanta can be realized by putting a $2\pi$ phase winding in one of the two superconducting components $\Delta_{\pm\bm{Q}}$ of the unidirectional PDW order parameter. Thus, we will require the order parameter $\Delta_{\bm{Q}}(\bm{r})$ to have a unit vortex and set the other component to be real $\Delta_{-\bm{Q}}(\bm{r})=|\Delta_{-\bm{Q}}(\bm{r})|$, and we set its phase $\theta_{-\bm{Q}}=0$
A half-vortex is an extremal solution of the PDW free energy $\mathcal{F}_{pdw}$ which at long distances behaves as a vortex in $\Delta_{\bm{Q}}$
@@
$\lim_{|\bm{r}|\to\infty}\Delta_{-\bm{Q}}(\bm{r})$ $=\Delta_{pdw}$ (3.1)
@@
On the other hand, since $\Delta_{\bm{Q}}$ has a vortex, its amplitude must vanish at the origin. Instead, $\Delta_{-\bm{Q}}$ does not have to vanish at the origin and it will take some finite value, which we denote as $\Delta_{ff}$. In other words, in the presence of a half-vortex the superconducting order has an FF component. Hence, at short distances the half-vortex must have the behavior
@@
with $\bm{n}^{2}(\bm{r})=1$ everywhere, and where $\Delta$ will be determined below. Here we used that $\Delta_{-\bm{Q}}$ is a real field. With these assumptions the free energy density of the PDW, Eq.(2.3), becomes (with $\bm{n}=(n_{x},n_{y},n_{z})$)
@@
$\bar{\Delta}=\sqrt{\frac{2(T_{c}^{pdw}-T)}{4u-|\gamma|}}$ (3.6)
@@
$\mathcal{F}_{nlsm}={\bar{\kappa}}(\nabla\bm{n})^{2}-vn_{z}^{2}(n_{x}^{2}+n_{y}^{2})+\textrm{const}.$ (3.7)
@@
We will now construct the half-vortex of the PDW using the non-linear sigma model (NLSM) of Eq.(3.7). We assume that the ordering wavevector of the PDW is oriented along the $x$ axis and write $\mathbf{Q}\equiv Q\mathbf{e}_{x}$. We will also assume that the phase field $\varphi(\mathbf{r})$ winds by $2\pi$, and we define its branch cut along the positive $x$-axis. With these assumptions we define the unit vector
$\mathbf{e}_{\mathbf{r}}=\cos(\varphi(\mathbf{r}))\mathbf{e}_{x}+\sin(\varphi(\mathbf{r}))\mathbf{e}_{y}$ (3.8)
@@
$\mathbf{n}(\mathbf{r})=\text{sin}\,(\alpha(\mathbf{r}))\mathbf{e}_{\mathbf{r}}+\cos(\alpha(\mathbf{r}))\mathbf{e}_{z}$ (3.9)
In section 2 we defined the half-vortex as a configuration in which the phase field of $\Delta_{\bm{Q}}$ winds by $2\pi$ at infinity while the order parameter field $\Delta_{-\bm{Q}}$ does not wind and is defined to be real. The order parameters of the PDW in the half-vortex state are required to obey the boundary conditions of Eq.(3.1) (at long distances) and Eq.(3.2) (at short distances). As a result, the half-vortex has an FF-type order within the core and asymptotically far from the core is of LO-type. Such a state breaks inversion symmetry in the core of the half-vortex. In terms of the NLSM field $\bm{n}(\bm{r})$ the boundary conditions of Eqs.(3.1) and (3.2) become
$\lim_{r\rightarrow 0}{\bm{n}}(\mathbf{r})$ $=(0,0,1)$ (3.10)
$\lim_{r\rightarrow\infty}{\bm{n}}(\bm{r})$ $=\frac{1}{\sqrt{2}}(\cos\varphi(\bm{r}),\sin\varphi(\bm{r}),1)$ (3.11)
@@
$\lim_{r\rightarrow 0}\alpha(r)=0,\;\;\;\lim_{r\rightarrow\infty}\alpha(r)=\pi/4$ (3.12)
@@
After performing some algebra, we can recast the total free energy of the NLSM, Eq.(3.7) into the following form
@@
$\frac{d^{2}\alpha}{dt^{2}}=\frac{1}{2}\Big{[}1-\text{cos}(2\alpha(t))e^{2t}\Big{]}\text{sin}(2\alpha(t))$ (3.15)
@@
$\lim_{t\to-\infty}\alpha(t)=0,\qquad\lim_{t\to\infty}\alpha(t)=\pi/4$ (3.16)
A numerical solution of the equation of motion of Eq.(3.15) yields the optimal solution for the half-vortex. Plots of the magnitudes of the PDW components (in units of $\bar{\Delta}$) as a function of distance from the vortex core are provided in figure 2a. As it is clear from this figure, as the amplitude $\Delta_{\bm{Q}}(r)$ decreases as $r\to 0$, the amplitude of $\Delta_{-{\bm{Q}}}(r)$ increases as $r\to 0$. In other words, the core of the half-vortex behaves as a FF state which breaks inversion symmetry. The parameter $r_{0}$ can be used to define the radius of the half-vortex and it is set at the inflection of the $\Delta_{-{\bm{Q}}}(\mathbf{r})$ field (at $r/r_{0}=1$ in figure 2), which is related to the coherence length of the Cooper pairs. We can then see that the degree of inversion symmetry breaking depends on the area of the core of the half-vortex. In the absence of the uniform component $\Delta_{0}$ the branch cut of the half-vortex is unobservable resulting in a free energy that is only logarithmically divergent. However, if $\Delta_{0}\neq 0$ the branch cut becomes observable and behaves as a domain wall. In this case the energy of the half-vortex becomes linearly divergent.
@@
$\Delta(\bm{r})=\bar{\Delta}f(r/r_{0})\exp(i\varphi(\bm{r}))$ (3.17)

### v4

@@ 2 Landau-Ginzburg Theory
In this section we review the Landau-Ginzburg (LG) theory for a unidirectional PDW state [25], which will be used to describe the profiles of our order parameters in the presence of static topological defects. In a a later section we will calculate the local density of states (LDOS) in the presence of topological defects that we will describe in this section. Because our defects are static, we will be neglecting fluctuations, so our analysis here will be done at the level of mean field theory. Furthermore, we also work deep in the PDW phase where the order parameters have well defined local amplitudes, which is justified at low temperatures.
@@ 2.1 Free energy of the PDW state
The free energy of the LG theory is chosen to have the form:
$\mathcal{F}[\Delta_{0},\Delta_{\mathbf{Q}},\Delta_{-\mathbf{Q}}]=\mathcal{F}_{\text{sc}}[\Delta_{0}]$ $+\mathcal{F}_{\text{pdw}}[\Delta_{\mathbf{Q}},\Delta_{-\mathbf{Q}}\big{]}.$ (2.1)
The first term in Eq.  (2.1) describes uniform superconductivity, which takes the familiar form
$\mathcal{F}_{\text{sc}}=\frac{1}{2m^{*}}\Big{|}\big{(}-i\nabla+2e\mathbf{A}\big{)}\Delta_{0}\Big{|}^{2}+\frac{a}{2}|\Delta_{0}|^{2}+\frac{b}{2}|\Delta_{0}|^{4}.$ (2.2)
@@
The second term in Eq.  (2.1) describes the free energy of the unidirectional PDW. In Eq. (1.1) we presented the expansion of the local pairing amplitude $\Delta(\bm{r})$ in its Fourier components $\Delta_{\pm\bm{Q}}(\bm{r})$, where $\Delta_{\pm\bm{Q}}(\bm{r})$ are the order parameters of the unidirectional PDW state with ordering wave vector $\bm{Q}$. Since the local pairing amplitude $\Delta(\bm{r})$ is a complex field, the two PDW components $\Delta_{\pm\bm{Q}}$ are not the complex conjugate of each other (as they would have been for a CDW) but are two independent complex fields. The PDW free energy has the form [25, 26, 45, 1]:
$\mathcal{F}_{\text{pdw}}[\Delta_{\mathbf{Q}},\Delta_{-\mathbf{Q}}\big{]}=$ $\kappa\Big{(}|\nabla\Delta_{\mathbf{Q}}|^{2}+|\nabla\Delta_{-\mathbf{Q}}|^{2}\Big{)}+r\Big{(}|\Delta_{\mathbf{Q}}|^{2}+|\Delta_{-\mathbf{Q}}|^{2}\Big{)}+u\Big{(}|\Delta_{\mathbf{Q}}|^{2}+|\Delta_{-\mathbf{Q}}|^{2}\Big{)}^{2}+\gamma|\Delta_{\bm{Q}}|^{2}|\Delta_{-\bm{Q}}|^{2},$ (2.3)
where $r=T-T_{c}^{\rm pdw}$, and $T_{c}^{\rm pdw}$ is the (mean-field) critical temperature for the PDW superconductor. We will assume that the coupling constant $\gamma<0$ as needed for an attractive interaction needed for a PDW, an LO-type state. Instead, a repulsive value of the coupling constant, $\gamma>0$, favors an FF type state.
@@
In many superconductors of interest, such as cuprates Bi2Sr2CaCu2O8+δ and La2-xBaxCuO4 in the Meissner state, a uniform superconducting order parameter $\Delta_{0}$ is present. This results in the following additional “lock-in” terms that couple the three superconducting order parameters and it must be added to the free energy of Eq. (2.1):
@@ 2.2 Order parameters of the PDW state
$\rho_{\bm{Q}}(\bm{r})$ $=\Delta_{0}(\bm{r})\Delta_{\bm{Q}}^{*}(\bm{r}),$ (2.5)
$\rho_{2\bm{Q}}(\bm{r})$ $=\Delta_{\bm{Q}}(\bm{r})\Delta_{-\bm{Q}}^{*}(\bm{r}),$ (2.6)
$\Delta_{\text{4e}}({\bm{r}})$ $=\Delta_{\mathbf{Q}}({\bm{r}})\Delta_{-\mathbf{Q}}({\bm{r}}).$ (2.7)
@@
Following the analysis of Berg et al. [25] we decompose the phase fields of the PDW order parameters as
$\theta_{\pm\bm{Q}}(\bm{r})=\theta_{+}(\bm{r})\pm\theta_{-}(\bm{r}).$ (2.8)
@@
$\Delta_{\pm\bm{Q}}(\bm{r})$ $\to\exp(i(\vartheta_{+}\pm\vartheta_{-}))\Delta_{\pm\bm{Q}}(\bm{r}),$ (2.9)
$\Delta_{0}(\bm{r})$ $\to\exp(i\vartheta_{0})\Delta_{0}(\bm{r}),$ (2.10)
$\rho_{\bm{Q}}(\bm{r})$ $\to\exp(i(\vartheta_{0}-\vartheta_{+}))\exp(i\vartheta_{-})\rho_{\bm{Q}}(\bm{r}),$ (2.11)
$\rho_{2\bm{Q}}(\bm{r})$ $\to\exp(i2\vartheta_{-})\rho_{2\bm{Q}}(\bm{r}),$ (2.12)
$\Delta_{4e}(\bm{r})$ $\to\exp(i2\vartheta_{+})\Delta_{4e}(\bm{r}),$ (2.13)
@@
$\rho(\bm{r})=\bar{\rho}+\rho_{\bm{K}}(\bm{r})\exp(i\bm{K}\cdot\bm{r})+\rho_{-\bm{K}}(\bm{r})\exp(-i\bm{K}\cdot\bm{r})+\ldots,$ (2.14)
@@
$\mathcal{F}_{PDW-CDW}=g\,\rho_{\bm{K}}\Delta_{\bm{Q}}\Delta^{*}_{-\bm{Q}}+{\rm c.c.},$ (2.15)
where $g$ is a coupling constant [25]. Translation invariance then requires that the mutual commensurability condition is satisfied. The same requirement exists for a coupling between a spin density wave $SDW$. This effect is seen in La2-xBaxCuO4 at the charge-ordering transition [50] and in La1.8-zEu0.2SrxCuO4 [51] at low temperatures. The existence of this interaction yields some interesting physics not allowed for an uniform SC state: the PDW is sensitive to charged impurities due to their coupling to the CDW. Thus, unlike uniform superconducting order, which is weakened only by disorder, even small amounts of charge disorder destroys true long-range incommensurate PDW order.
@@ 2.3 Topological defects of the PDW state
$q_{v}=\frac{1}{2}(m_{\bm{Q}}+m_{-\bm{Q}}),\qquad q_{d}=\frac{1}{2}(m_{\bm{Q}}-m_{-\bm{Q}}).$ (2.16)
We will label the topological defects by the combinations $(q_{v},q_{d})$, the vorticity and dislocation charges, or equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})$. In the simplest cases they are (1) the superconducting (Abrikosov) vortex with topological charges $(q_{v},q_{d})=(\pm 1,0)$ [or, equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})=(\pm 1,\pm 1)$], (2) the half-vortex (bound to a single CDW dislocation) with topological charges $(q_{v},q_{d})=(\pm 1/2,\pm 1/2)$ [or, equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})=(\pm 1,0)$], and (3) the CDW double dislocation with topological charges $(q_{v},q_{d})=(0,\mp 1)$ [or, equivalently, $(m_{\bm{Q}},m_{-\bm{Q}})=(\pm 1,\mp 1)$]. These identifications imply that a conventional superconducting vortex is equivalent to both PDW order parameters $\Delta_{\pm\bm{Q}}$ having a vortex. The half-vortex is equivalent to a vortex in $\Delta_{\bm{Q}}$ but not in $\Delta_{-\bm{Q}}$ (and vice versa), and it has a single-dislocation, as required by Eq.(2.6). Finally, a double dislocation is a vortex in $\Delta_{\bm{Q}}$ and an anti-vortex in $\Delta_{-\bm{Q}}$ (and vice versa).
@@
In this paper we will focus primarily on the properties of the half-vortex which we will regard as a static topological defect of an ordered PDW state which we will take to be of the Larkin-Ovchinnikov (LO) type [23]. Hence, we will assume that at long distances the PDW amplitudes are equal and constant in space, $|\Delta_{\bm{Q}}|=|\Delta_{-\bm{Q}}|$. However, at the core of the half-vortex one of these two amplitudes, say $\Delta_{\bm{Q}}$, must vanish while the other amplitude does not. As a result, the core of the half-vortex is in a Fulde-Ferrell (FF) state [24], and inversion symmetry is broken at the core of the half-vortex since $|\Delta_{\bm{Q}}|\neq|\Delta_{-\bm{Q}}|$. On the other hand, none of these considerations apply to either the Abrikosov vortex or to the double dislocation.
As in all superconductors, the Abrikosov vortex of the PDW arises in the presence of a magnetic field. On the other hand, the half-vortex and the double dislocation can appear due to the interaction of the PDW state with sufficiently strong charged impurities. This is possible since the CDW order parameter $\rho_{2\bm{Q}}(\bm{r})$ of the PDW has a linear coupling to charged impurities potentials, whereas gauge invariance requires that the superconducting order parameters interact only quadratically through $|\Delta_{\pm\bm{Q}}(\bm{r})|^{2}$.
The properties of all three topological defects will be discussed in Sec. 4, where we specify the form of the order parameters used in the numerics. The associated charge density wave patterns of each defect will be compared in Sec. 4.2 as well as their spectral functions in Sec. 4.5.
An extremely useful experimental technique for detecting and visualizing CDW patterns is scanning tunneling microscopy (STM) [52, 53, 54, 55]. STM has been used to investigate in great detail the charge order present in the superconducting phase of Bi2Sr2CaCu2O8+δ [56, 57, 58] and in the vortex halo [15, 16]. Relevant to the existence of PDW order is the experimental evidence for static (pinned) half-vortices in the superconducting phase of Bi2Sr2CaCu2O8+δ found in STM experiments by Du et al. [14]. These authors argued that the jumps in the PDW SC winding number are located around the charge dislocations. In this regime the term proportional to $\beta_{2}$ in Eq. (2.4) locks the phase of the PDW order parameter to the phase of the uniform superconducting state. The result is a confinement of the half-vortices into (half) vortex-anti-vortex pairs. Hence, in the phase in which the uniform order parameter is present, $\Delta_{0}\neq 0$, the energy of the half-vortex is linearly divergent and half-vortices cannot exist in isolation. In this regime charged impurity potentials can separate the half-vortices and anti-half-vortices as static defects.
@@ 3 The Half-Vortex
In this section we will model the profile of our half-vortex of an unidirectional PDW and we will assume that there is no uniform SC order. In Sec. 4 we will embed the configuration of the half-vortex in a Bogoliubov-de Gennes Hamiltonian for the $d$-wave superconducting state of a CuO2 plane and investigate its effects on the electronic structure. The Abrikosov vortex and the double dislocation will also be considered in later sections, so we mention their solutions here as well. We should note that the associated CDW dislocation was given in Ref. [45].
We will seek an extremal solution of the Landau-Ginzburg free energy density $\mathcal{F}_{pdw}(\Delta_{\bm{Q}}(\bm{r}),\Delta_{-\bm{Q}}(\bm{r}))$ for the PDW order given in Eq.(2.3). In Sec. 2 we showed that a vortex with half magnetic flux quanta can be realized by putting a $2\pi$ phase winding in one of the two superconducting components $\Delta_{\pm\bm{Q}}$ of the unidirectional PDW order parameter. Thus, we will require the order parameter $\Delta_{\bm{Q}}(\bm{r})$ to have a unit vortex and set the other component to be real $\Delta_{-\bm{Q}}(\bm{r})=|\Delta_{-\bm{Q}}(\bm{r})|$, and we set its phase $\theta_{-\bm{Q}}=0$
A half-vortex is an extremal solution of the PDW free energy $\mathcal{F}_{pdw}$ which at long distances behaves as a vortex in $\Delta_{\bm{Q}}$:
@@
$\lim_{|\bm{r}|\to\infty}\Delta_{-\bm{Q}}(\bm{r})$ $=\Delta_{pdw},$ (3.1)
@@
On the other hand, since $\Delta_{\bm{Q}}$ has a vortex, its amplitude must vanish at the origin. Instead, $\Delta_{-\bm{Q}}$ does not have to vanish at the origin, and it will take some finite value, which we denote as $\Delta_{ff}$. In other words, in the presence of a half-vortex the superconducting order has an FF component. Hence, at short distances the half-vortex must have the behavior
@@
with $\bm{n}^{2}(\bm{r})=1$ everywhere, and where $\Delta$ will be determined below. Here we used that $\Delta_{-\bm{Q}}$ is a real field. With these assumptions the free energy density of the PDW, Eq. (2.3), becomes [with $\bm{n}=(n_{x},n_{y},n_{z})$]
@@
$\bar{\Delta}=\sqrt{\frac{2(T_{c}^{pdw}-T)}{4u-|\gamma|}}.$ (3.6)
@@
$\mathcal{F}_{nlsm}={\bar{\kappa}}(\nabla\bm{n})^{2}-vn_{z}^{2}(n_{x}^{2}+n_{y}^{2})+\textrm{const},$ (3.7)
@@
We will now construct the half-vortex of the PDW using the non-linear sigma model (NLSM) of Eq.(3.7). We assume that the ordering wave vector of the PDW is oriented along the $x$ axis and write $\mathbf{Q}\equiv Q\mathbf{e}_{x}$. We will also assume that the phase field $\varphi(\mathbf{r})$ winds by $2\pi$, and we define its branch cut along the positive $x$ axis. With these assumptions we define the unit vector
$\mathbf{e}_{\mathbf{r}}=\cos[\varphi(\mathbf{r})]\mathbf{e}_{x}+\sin[\varphi(\mathbf{r})]\mathbf{e}_{y},$ (3.8)
@@
$\mathbf{n}(\mathbf{r})=\text{sin}[\alpha(\mathbf{r})]\mathbf{e}_{\mathbf{r}}+\cos[\alpha(\mathbf{r})]\mathbf{e}_{z}.$ (3.9)
In Sec. 2 we defined the half-vortex as a configuration in which the phase field of $\Delta_{\bm{Q}}$ winds by $2\pi$ at infinity while the order parameter field $\Delta_{-\bm{Q}}$ does not wind and is defined to be real. The order parameters of the PDW in the half-vortex state are required to obey the boundary conditions of Eq. (3.1) (at long distances) and Eq. (3.2) (at short distances). As a result, the half-vortex has an FF type order within the core and asymptotically far from the core is of LO-type. Such a state breaks inversion symmetry in the core of the half-vortex. In terms of the NLSM field $\bm{n}(\bm{r})$ the boundary conditions of Eqs. (3.1) and (3.2) become
$\lim_{r\rightarrow 0}{\bm{n}}(\mathbf{r})$ $=(0,0,1),$ (3.10)
$\lim_{r\rightarrow\infty}{\bm{n}}(\bm{r})$ $=\frac{1}{\sqrt{2}}(\cos\varphi(\bm{r}),\sin\varphi(\bm{r}),1),$ (3.11)
@@
$\lim_{r\rightarrow 0}\alpha(r)=0,\;\;\;\lim_{r\rightarrow\infty}\alpha(r)=\pi/4.$ (3.12)
@@
After performing some algebra, we can recast the total free energy of the NLSM, Eq.(3.7) into the following form:
@@
$\frac{d^{2}\alpha}{dt^{2}}=\frac{1}{2}\Big{[}1-\text{cos}(2\alpha(t))e^{2t}\Big{]}\text{sin}[2\alpha(t)]$ (3.15)
@@
$\lim_{t\to-\infty}\alpha(t)=0,\qquad\lim_{t\to\infty}\alpha(t)=\pi/4.$ (3.16)
A numerical solution of the equation of motion (3.15) yields the optimal solution for the half-vortex. Plots of the magnitudes of the PDW components (in units of $\bar{\Delta}$) as a function of distance from the vortex core are provided in Fig. 2a. As is clear from this figure, as the amplitude $\Delta_{\bm{Q}}(r)$ decreases as $r\to 0$, the amplitude of $\Delta_{-{\bm{Q}}}(r)$ increases as $r\to 0$. In other words, the core of the half-vortex behaves as a FF state which breaks inversion symmetry. The parameter $r_{0}$ can be used to define the radius of the half-vortex and it is set at the inflection of the $\Delta_{-{\bm{Q}}}(\mathbf{r})$ field (at $r/r_{0}=1$ in Fig. 2), which is related to the coherence length of the Cooper pairs. We can then see that the degree of inversion symmetry breaking depends on the area of the core of the half-vortex. In the absence of the uniform component $\Delta_{0}$ the branch cut of the half-vortex is unobservable resulting in a free energy that is only logarithmically divergent. However, if $\Delta_{0}\neq 0$ the branch cut becomes observable and behaves as a domain wall. In this case the energy of the half-vortex becomes linearly divergent.
@@
$\Delta(\bm{r})=\bar{\Delta}f(r/r_{0})\exp(i\varphi(\bm{r})),$ (3.17)